### v1

@@ 6.4 Criteria-based evaluation
Figure 11 shows the remarkable performance of the Large Language Model’s (LLM) in the evaluated metrics, demonstrating consistency with slight variations in the scores. On average, the answers provided are accurate and in line with the reference, though there might be room for improvement. While responses generally provide useful information, there are instances where additional depth could enhance their utility. In addition, most answers usually include a remarkable balance between conciseness and effectiveness, showing a proficient level of brevity, as well as clear and logical organization. The overall results show no notable differences in these metrics across the three scenarios. This suggests that messages within the Rosbag related to the introduction of obstacles or cancellation of goals do not notably influence the general criteria evaluation outcomes.
@@
Figure 12: Large Language Model criteria-based evaluation by categories. Comparative visualizations for: (a) Scenario 1, (b) Scenario 2, and (c) Scenario 3.
Concerning general questions about about events and the progression of the navigation task, we observe variations in scores across Scenarios 2 and 3. While coherence and conciseness are even improved, Scenario 1, involving direct navigation to three goals, outperforms in accuracy and helpfulness. This fact suggests that effective communication on general queries benefits from the absence of obstacles or goal cancellations, enhancing overall understanding.
In the category ’Trajectory Planning and Replanning’, the questions are effectively addressed in all the three scenarios, showing scores consistently above 8 in every evaluated metric, even when obstacles are introduced or a goal is cancelled. These results underlines the system’s proficiency in handling trajectory-related questions despite potential new challenges.
@@
Concerning the ’Specifics about Goals and Locations’ category, the system consistently provide valuable goal and location information, even in the presence of obstacles or goal cancellations. This suggests a robust capacity to understand and localize specific details regardless of challenges introduced in each scenario.
@@ 7 Conclusions
This work presents an accountability and explicability architecture, implemented for ROS-based mobile robots. Our approach comprises a black box component for accountability that effectively addresses real-time recording challenges, ensuring message integrity and verifiability through blockchain technology, and an explainability component. This last component is based on the use of RAG and LLM through a private question-answering system designed to use the information recorded by the first component as context. This solution demonstrates the feasibility of obtaining accurate, coherent, concise and helpful explanations for robot actions from raw data stored in Rosbag files, thereby enhancing human-robot interaction. Additionally, both main components have been integrated through the development of a third main functionality in charge of filtering non-relevant information and making messages closer to the final user, due to their specific nature can introduce challenges in their comprehension process. Furthermore, this last component releases software developers from making direct calls to the logging API, easing the development process and improving the scalability and adaptability of our solution for every navigation task performed by a ROS-based robot.
@@
Additionally, our RAG approach could improve in terms of retrieving the most relevant chunk of information from the context. This enhancement may involve exploring the use of separate collections within the vector database or employing a hybrid approach, such as integrating knowledge graphs (KG) with vector indexes. These approaches can help to improve the accuracy, quality and overall consistency of the answers obtained across several scenarios, thereby enhancing the performance of our proposed system.
@@
Data and tools are available in a public repository.

### v2

@@ 6.4 Criteria-based Evaluation
Figure 11 shows the remarkable performance of the selected LLM in the evaluated metrics, demonstrating consistency with slight variations in the scores. On average, the answers provided are accurate and in line with the reference, though there might be room for improvement. While responses generally provide useful information, there are instances where additional depth could enhance their utility. In addition, most answers usually include a remarkable balance between conciseness and effectiveness, showing a proficient level of brevity, as well as clear and logical organization. The overall results show no notable differences in these metrics across the three scenarios. This suggests that messages within the Rosbag related to the introduction of obstacles or cancellation of goals do not notably influence the general criteria evaluation outcomes.
@@
Figure 12: LLM criteria-based evaluation by categories. Comparative visualizations for: (a) Scenario 1, (b) Scenario 2, and (c) Scenario 3.
Concerning general questions about events and the progression of the navigation task, we observe variations in scores across Scenarios 2 and 3. While coherence and conciseness are even improved, Scenario 1, involving direct navigation to three goals, outperforms in accuracy and helpfulness. This fact suggests that effective communication on general queries benefits from the absence of obstacles or goal cancellations, enhancing overall understanding.
In the category ’Trajectory Planning and Re-planning’, the questions are effectively addressed in all three scenarios, showing scores consistently above 8 in every evaluated metric, even when obstacles are introduced or a goal is cancelled. These results underline the system’s proficiency in handling trajectory-related questions despite potential new challenges.
@@
Concerning the ’Specifics about Goals and Locations’ category, the system consistently provides valuable goal and location information, even in the presence of obstacles or goal cancellations. This suggests a robust capacity to understand and localize specific details regardless of the challenges introduced in each scenario.
@@ 7 Conclusions
This work presents an accountability and explicability architecture, implemented for Robot Operating System (ROS)-based mobile robots. Our approach comprises a black box component for accountability that effectively addresses real-time recording challenges, ensuring message integrity and verifiability through blockchain technology, and an explainability component. This last component is based on the use of Retrieval Augmented Generation (RAG) and Large Language Models (LLMs) through a private question-answering system designed to use the information recorded by the first component as context. This solution demonstrates the feasibility of obtaining accurate, coherent, concise and helpful explanations for robot actions from raw data stored in Rosbag files, thereby enhancing human-robot interaction. Additionally, both main components have been integrated through the development of a third main functionality in charge of filtering non-relevant information and making messages closer to the final user, due to their specific nature can introduce challenges in their comprehension process. Furthermore, this last component releases software developers from making direct calls to the logging Application Programming Interface (API), easing the development process and improving the scalability and adaptability of our solution for every navigation task performed by a ROS-based robot.
@@
Additionally, our RAG approach could improve in terms of retrieving the most relevant chunk of information from the context. This enhancement may involve exploring the use of separate collections within the vector database or employing a hybrid approach, such as integrating Knowledge Graphs with vector indexes. Furthermore, by incorporating real-time explanation generation mechanisms into our solution, the responsiveness and adaptability of the system could be improved. These approaches can help to improve the accuracy, quality and overall consistency of the answers obtained across several scenarios, thereby enhancing the performance of our proposed system.
Appendix A Supplementary Materials
A.1 Prompt Template
[…]
Figure 13: Prompt template used to generate explanations with Llama 2 LLM.
A.2 Explanations Obtained
The 288 questions and answers obtained through the explanation generation and evaluation process are available on the aforementioned mentioned Github repository  [^4] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
gpt4_eval_results/LLMs_QA_Results.xlsx
. Additionally, a Python notebook has been developed to interact with the content of the previous sheet. This notebook is available on  [^5] https://github.com/laurafbec/immutable_explainable_BBR/blob/main/
gpt4_eval_results/GPT4 Evaluation Results.ipynb
@@ A.2 Explanations Obtained
Data and tools are available in a public repository  [^6] https://github.com/laurafbec/immutable_explainable_BBR.